### 03 Mar 2024 v1

# Towards Comprehensive Vietnamese Retrieval-Augmented Generation and Large Language Models

Abstract
This paper presents our contributions towards advancing the state of Vietnamese language understanding and generation through the development and dissemination of open datasets and pre-trained models for Vietnamese Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs).
Keywords RAG, LLMs, Open Datasets
1 Introduction
We hope that the research community, both in Vietnam and around the world, will join forces in the endeavor to construct large and high-quality datasets for the training and evaluation of Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs) for Vietnamese. By collaborating on this front, we can collectively push the boundaries of what’s possible in natural language processing for Vietnamese, unlocking new opportunities for innovation and application in the field.
Together, let’s work towards an open scientific community that benefits everyone!
2 Main contributions
Our main contributions are as follows:
  • A massive Vietnamese NewsCorpus dataset of around 32M articles, a substantial 53 GB in size, rigorously cleaned, deduplicated, and formatted specifically for the continual pre-training of LLMs.
  • An extensive Vietnamese NewsSapo dataset, structured in a “title-abstract-contents” format, specifically designed to enhance the training of sentence/passage embeddings.
  • An additional large-scale Vietnamese NewsCategory dataset in a “text-category” format, specifically designed for text classification task.
  • Vietnamse Alpaca datasets, tailored for supervised fine-tuning LLMs.
  • Synthetic self-chat and roleplay realm datasets are developed for enhancing the conversation capability of LLMs through supervised fine-tuning.
  • A good Vietnamese bi-encoder model is presented for advanced sentence embedding tasks.
  • We also offer two base models, Vietnamese LLaMA2-7b, which have been further pretrained on an expansive corpus of Vietnamese text, 40 GB and 120 GB respectively, derived from LLaMA2, marking a significant advancement in the understanding and generation of the Vietnamese language.
3 Details
3.1 Vietnamese NewsCorpus dataset
The Binhvq News Corpus [1], a widely used dataset featuring approximately 20 million articles from diverse sources, received its last update in May 2021. To enhance this collection, we gathered an additional 10 million articles up until November 2023. By integrating these newly acquired articles with the existing Binhvq News Corpus, we have created an extensive Vietnamese News Corpus comprising about 32M articles. Subsequent fuzzy deduplication was conducted to remove duplicate articles, resulting in 53 GB of clean data, which is ready for the continual pretraining of LLMs.
3.2 Vietnamese NewsSapo dataset
The Vietnamese NewsSapo dataset was constructed to training of sentence/passage embeddings. Our dataset is structured in a ”title-abstract-contents” format, where each news article is represented by a tuple of (title, abstract, content). The content is the main text body of the article and has been processed to remove images, videos, and other non-textual elements. The dataset contains 31,728,183 triples.
3.3 Vietnamese NewsCategory dataset
The Vietnamese NewsCategory dataset is constructed similarly to the THUCTC dataset [2]. The dataset is collected from VnExpress and is extracted for classification tasks. It contains 596,524 samples, each of which consists of five fields: id (index), title, sapo (summary), content (the contents of articles), label (article topic).
[…]
3.4 Vietnamese Alpaca datasets
3.4.1 Standard Vietnamese Alpaca
This dataset is specifically tailored for Vietnamese, drawing inspiration from the methodologies of Stanford Alpaca [3] and Self-Instruct [4]. The construction of this dataset involved a systematic two-step process:
  • Creation of Vietnamese Seed Tasks: Following the idea in Self-Instruct [4], we carefully developed a wide-ranging collection of seed tasks for Vietnamese. This was achieved using GPT-4, alongside meticulous manual crafting.
  • Generation of Instructions: Leveraging the seed tasks prepared in the first step, we engaged in the instruction generation technique inspired by Stanford Alpaca [3]. Utilizing GPT-4, GPT-3.5 turbo, and GPT-3.5-instruct, we generated 50K instructions. This process was carried out using various configurations to ensure the production of a rich and diverse array of linguistic scenarios.
3.4.2 Modified Vietnamese Alpaca
This dataset has been developed using a similar method to that of the Standard Vietnamese Alpaca, with a notable difference in format. While the Standard Vietnamese Alpaca adopts the “instructions/inputs/outputs” format as suggested by [3], our method, influenced by [5], utilizes the “input/output” format. This alteration allows for the generation of more diverse samples and more lengthy output, employing GPT-4, GPT-3.5, and GPT-3.5-instruct. The resulting Modified Vietnamese Alpaca dataset contains 50K samples.
3.5 Vietnamese Self-chat dataset
This dataset is designed to enhance the model’s ability to engage in multi-turn conversations with humans. We follow two steps:
  • Instruction Generation: We employ the methodology outlined in Self-Instruct [4] to craft a diverse set of instructions.
  • Synthetic Self-Chat Conversations: Building upon the instructions generated in the first step, we draw inspiration from Baize [6] to generate synthetic multi-turn dialogues. These simulations serve as practical scenarios for the model to learn from and adapt to dynamic conversation flows.
By combining these two steps, we aim to create a robust and versatile dataset that empowers the model to navigate and contribute effectively in complex conversational scenarios. This dataset serves as a valuable resource for refining the model’s language understanding and response generation capabilities in the context of human-like dialogue.
3.6 Vietnamese Roleplay Realm dataset
This is a dataset of GPT-generated characters made to increase the ability of open-source language models to role-play. It contains 446 characters generated by GPT-3.5. To construct this dataset, we follow four steps:
  • Character Generation: Creates a set of fictional characters with GPT-3.5 based on a prompt and a seed list of characters. The generated output fields for each character are “name”, “context”, “greeting”, and “example_dialogue”.
  • Topic Generation: We then created conversation topics for each character, drawing from their descriptions. The output field for this step is ”topics”. We generates 20 topics for each character.
  • Dialogue generation: Based on the character descriptions and topics, we generated dialogues. The output for this step is encapsulated in the ”dialogues” field.
  • Checking and Refining: Given that the dataset may contain errors in Vietnamese, a review and correction process is necessary to ensure accuracy and refinement.
3.7 Vietnamese Bi-encoder
This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space and can be used for tasks like clustering or semantic search.
We train the model on a merged training dataset that consists of:
  • MS Macro (translated into Vietnamese)
  • SQuAD v2 (translated into Vietnamese)
  • 80% of the training set from the Legal Text Retrieval Zalo 2021 challenge
We use phobert-base-v2 [7] as the pre-trained backbone.
Here are the results on the remaining 20% of the training set from the Legal Text Retrieval Zalo 2021 challenge:
Table 1: Results on the Legal Text Retrieval Zalo 2021 challenge
Pretrained Model Training Datasets Acc@1 Acc@10 Acc@100 Pre@10 MRR@10
Vietnamese-SBERT - 32.34 52.97 89.84 7.05 45.30
PhoBERT-base-v2 MSMACRO 47.81 77.19 92.34 7.72 58.37
PhoBERT-base-v2 MSMACRO + SQuADv2.0 + 80% Zalo 73.28 93.59 98.85 9.36 80.73
3.8 Vietnamese LLaMA2
3.8.1 Vietnamese LLaMA2-7b-40Gb
We employed SentencePiece to retrain a Vietnamese tokenizer with a vocabulary size of 20K. No Vietnamese word segmentation was used. We then merged this vocabulary with the original one of LLaMA2, removing duplicate tokens. The new tokenizer significantly improves when encoding Vietnamese text, reducing the number of tokens by 50% compared to ChatGPT and approximately 70% compared to the original LLaMA2.
We conducted a single-epoch continual pretraining, also known as incremental pretraining, using the LLaMA2-chat 7B model on a mixed dataset totaling 40.5 GB, comprised of:
  • 19 GB NewsCorpus [1]
  • 1.1 GB Vietnamese Wikipedia
  • 4.5 GB Vietnamese legal documents (crawled from thuvienphapluat and processed by ourselves)
  • 2.1 GB Vietnamese legal text (from C4-vi)
  • 1.1 GB English Books (sub-sampled from pg19)
  • 1.1 GB English Wikipedia (sub-sampled from 20220301.en wikipedia)
  • 10 GB English Text (sub-sampled from C4-en)
We trained the model on a DGX A100 system, utilizing four GPU A100 in 10 days (about 1000 GPU hours).
We also provide the LoRA part so that you can integrate it with the original LLaMA2-chat-7b by yourself.
3.8.2 Vietnamese LLaMA2-7b-120GB
Tokenizer
We enhance our previous tokenizer in Vietnamese-LLaMA2-7b-40GB by training SentencePiece on a more extensive collection of clean Vietnamese documents spanning diverse domains such as news, books, stock, finance, and laws. In contrast to the previous version, we follow the original LLaMA-2 paper to split all numbers into individual digits. Again, the updated tokenizer markedly enhances the encoding of Vietnamese text, cutting down the number of tokens by 50% compared to ChatGPT and approximately 70% compared to the original LLaMA2.
Pretraining data
Here are our data sources:
  • Vietnamese NewsCorpus described in Section 3.1
  • 1.3 GB Vietnamese Wikipedia (updated to October 2023)
  • 4.8 GB Vietnamese legal documents (clean and dedup)
  • 1.6 GB stock news (clean and dedup)
  • 43 GB Vietnamese text (subsampled from Culturax-vi [8])
  • 2.3 GB English Books (sub-sampled from pg19)
  • 2.2 GB English Wikipedia
  • 16 GB English text (subsampled from Culturax-en [8])
We then merge all data sources and perform the last deduplication, resulting in a final pretraining dataset of 124 GB, including 104 GB of Vietnamese text and 20 GB of English text.
Continual pretraining
We conduct a single-epoch continual pretraining using the LLaMA2-7B model. We trained the model on a DGX A100 system, utilizing four GPU A100 in 40 days (about 4000 GPU hours).
Hyperparameters are set as the same as the Vietnamese-LLaMA2-7b-40GB.
We also provide the LoRA part so that you can integrate it with the original LLaMA2-7b by yourself.
Training loss
The red line indicates the learning curve of Vietnamese-LLaMA2-7b-40GB, while the cyan one corresponds to the new model of 120 GB.
4 Conclusion
By opening access to our datasets and models, we extend an invitation to the broader research community to collaborate with us on the path to developing more inclusive, efficient, and accessible Vietnamese Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs).
Together, we can drive innovation, enhance linguistic inclusivity, and foster a rich ecosystem of NLP tools and technologies that bring substantial benefits to Vietnam.
5 Acknowledgments
This work was funded by NAVER Corporation.